The perimeter of a rectangle is 34ft. the length is 1ft. longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let X be the length and y be the width.
First equation 2x+2y= ?
write the second equation
x=y + ?
What is the length?
What is the width?

Answer by Cromlix(4381)

Width = x
Length = x + 1
Perimeter = 2* length + 2* width
34 = 2* (x + 1) + 2*x
34 = 2x + 2 + 2x
4x = 34 - 2
4x = 32
x = 8
Width = 8ft
Length = 9ft